ABSOLUTE INTENSITIES IN ACETYLENE: THE 7.5μm REGION

M. EL AZZAOUI, and \underline{J. VANDER AUWERA}, Laboratoire de Chimie Physique Moléculaire C. P. 160/09, Université Libre de Bruxelles, 50 avenue F. D. Roosevelt, B-1050 Brussels, Belgium.

Spectra of C_{2}H_{2} have been recorded at room temperature in the range from 1140 to 1490 cm^{−1} using a Bruker IFS120HR spectrometer. The interferograms were collected with a maximum optical path difference of 300 cm. Sample pressure/absorption path length products ranging from 1.3 to 330 hPa*cm were used. Absolute line intensities have been measured using direct non-linear least squares fitting of Voigt profiles — convoluted with an instrument line shape function — to the observed spectral data. Measurements have been performed in the ν_{4} + ν_{5} cold band and in hot bands originating from levels involving the bending modes ν_{4} and ν_{5} ^{a}. The results and analysis of the measurements will be presented and discussed.

aY. Kabbadj, M. Herman, G. Di Lonardo, L. Fusina, and J. W. C. Johns, J. Mol. Spectrosc. 150, 535 (1991)